### v2

# Benchmarking VLMs’ Reasoning About Persuasive Atypical Images

Abstract
Vision language models (VLMs) have shown strong zero-shot generalization across various tasks, especially when integrated with large language models (LLMs). However, their ability to comprehend rhetorical and persuasive visual media, such as advertisements, remains understudied. Ads often employ atypical imagery, using surprising object juxtapositions to convey shared properties. For example, Fig. 1 (e) shows a beer with a feather-like texture. This requires advanced reasoning to deduce that this atypical representation signifies the beer’s lightness.
We introduce three novel tasks, Multi-label Atypicality Classification, Atypicality Statement Retrieval, and Atypical Object Recognition, to benchmark VLMs’ understanding of atypicality in persuasive images. We evaluate how well VLMs use atypicality to infer an ad’s message and test their reasoning abilities by employing semantically challenging negatives. Finally, we pioneer atypicality-aware verbalization by extracting comprehensive image descriptions sensitive to atypical elements.
Our findings reveal that: (1) VLMs lack advanced reasoning capabilities compared to LLMs; (2) simple, effective strategies can extract atypicality-aware information, leading to comprehensive image verbalization; (3) atypicality aids persuasive advertisement understanding. Code and data are available at github.com/sinamalakouti/PersuasiveAdVLMBenchmark
11footnotetext: These authors contributed equally to this work.
1 Introduction
In visual media, particularly advertisements, creators employ creative visual rhetoric to capture attention and convey memorable, powerful messages. They intentionally deviate from realism, depicting objects in unique and atypical ways [34, 29]. Creative ads that are “out of the ordinary” or “connect objects that are usually unrelated” can generate twice as much revenue as non-creative ads [34].
Atypical imagery in ads often involves transforming objects metaphorically [43, 46]. These creative transformations are not random; they are carefully chosen to convey specific ideas  [43]. For example, Fig. 1(a) depicts a text box as tape to suggest silencing, while in (d), potato chips are shown as flames to metaphorically represent spiciness, borrowing properties from fire (hotness symbolizing flavor). Understanding these atypical images requires more than just recognizing objects. It requires advanced reasoning skills, including knowledge of cultural contexts and social norms, posing a significant challenge for AI systems.
Figure 1: Atypicality categories. We study four types of atypicality from [46]: Texture Replacement 1, Texture Replacement 2, Object Inside Objects, Object Replacement (defined in Sec. 3.1).
Modern pretrained vision-language models (VLMs), such as LLaVA [27, 26], demonstrate strong visual understanding across various tasks such as recognition [25], and capabilities like zero-shot generalizability. However, there is a lack of in-depth study on VLMs’ ability to understand complex persuasive images, such as advertisements.
We address this gap by introducing three novel tasks over PittAds [18] to evaluate VLMs’ understanding of atypicality: (1) multi-label atypicality classification, where the model predicts the type of atypicalities in the image; (2) atypicality statement retrieval, where the model retrieves correct atypicality statements describing the atypicality relation among objects; (3) atypical object recognition, where the model generates objects to complete an atypicality statement based on a given relation. These tasks are essential as prior works’ binary classification oversimplifies atypicality’s nuanced nature. Our evaluation shows that although VLMs struggle with direct atypicality inference, they can extract valuable information about atypical aspects.
Next, we investigate how atypicality influences understanding an ad’s message. We use the action-reason retrieval (ARR) task [18, 46], which requires models to identify the suggested action (e.g., “buy these chips") and its rationale (e.g., “because they are spicy"). However, to rigorously test the model’s reasoning, we introduce semantically challenging negative options, rather than mining hard negatives from other images [45, 19]. For example, we generate statements that include wrong action (e.g., “don’t buy these chips") or wrong rationale (e.g., “because they are sweet"). This prevents VLMs from ruling out negatives by merely comparing objects in the image and options. Our evaluation shows a significant performance drop when VLM is faced with hard negatives (e.g., LLaVA drops by 67.51%).
Finally, we hypothesize that deep atypicality understanding improves action-reason retrieval performance. We test this by proposing an atypicality-aware verbalization. We explore a few simple prompting strategies to construct a comprehensive, atypicality-sensitive ad verbalization, which is used to predict the atypicality statement. This statement, combined with the atypicality-aware verbalization, serves as input to an LLM classifier. This LLM integrates all the information to retrieve the final action-reason, effectively leveraging both visual understanding and high-level reasoning capabilities.
Our proposed framework achieves state-of-the-art performance on the ARR task. Interestingly, when a VLM is given both the image and our atypicality-aware verbalization, its performance on the ARR task actually declines (e.g., LLaVA($I+\mathcal{T}_{\mathcal{V}}$) shows a 1.71 point drop in performance compared to LLaVA($I$)) and it is significantly underperformed compared to LLM. This stark contrast highlights a critical gap: VLMs lack the advanced reasoning capabilities of LLMs when interpreting complex, atypical visual media. To summarize, our contributions are:
  1. 1.
    We introduce three novel tasks for understanding atypicality in persuasive media.
  2. 2.
    We pioneer the use of atypicality inference in action-reason retrieval and are the first to benchmark VLMs/LLMs for advertisement understanding.
  3. 3.
    We generate semantically challenging negatives using GPT-4 for action-reason retrieval, revealing VLMs’ reasoning limitations in interpreting atypical ads.
We hope this work inspires the inclusion of persuasive ads in VLM benchmarks, foster development of robust models for complex visual media, and offer insights for creating more effective advertisements.
2 Related Works
Creativity in advertising has long been of interest in advertising research. It has been broken down into categories, and its impact on the effectiveness of ads has been measured. Both [34] and [35] define the categories as originality, flexibility, synthesis, elaboration, and artistic value, which capture different shades of divergence from the ordinary. Atypicality most directly maps to synthesis. However, these creativity strategies have not been explored in computer vision for predicting the message of an ad.
Advertisement image understanding. The PittAds Dataset [18] introduced the action-reason retrieval task, establishing a baseline for automatic ad understanding. However, most studies have not explicitly captured advertising-specific strategies for this task, nor have they addressed atypicality. [45] incorporated symbolism, but the gains were minimal. Others utilized scene-text [14, 20], graph-based methods to incorporate external knowledge [47], and CLIP [33] for brand name understanding [19]. [5] used Automatic Speech Recognition, OCR, WikiData and BLIP-2 [24] to describe the stories of video ads. [2] analyzed metaphors in ads. Yet, the impact of atypicality on ad image understanding remains unexplored. The only exception is [17], which proposed a self-supervised approach to classify images as typical or atypical but did not classify the type of atypicality nor use them for action-reason prediction.
Vision-language models. We benchmark pre-trained VLMs and LLMs on tasks involving atypicality and advertisement image understanding, focusing on their zero-shot reasoning capabilities. We also use pre-trained VLMs to verbalize advertisement images for LLMs. Given the substantial computational power required for training and fine-tuning large models, off-the-shelf, frozen models are typically used [39, 3]. Techniques to align visual and textual features without parameter updates include optimizing image encoders [39], inserting cross-attention layers [3], prompt learning [55, 54, 9, 21], and employing external transformers [32, 24, 13, 26]. However, direct application of models like [24, 33] may miss hidden messages by focusing more on visual context than semantics. Note that we restrict our experiments to the zero-shot inference setting.
Language models for multi-modal reasoning. Recent studies [15, 51, 5, 44] have explored LLMs for reasoning tasks, including chain-of-thought reasoning [42]. Some works leverage LLMs in multi-modal reasoning. [51, 41, 52] extended chain-of-thought to a multimodal context.[23] uses an image-captioning model followed by ‘reasoning questions’ to aid an LLM in answering the main question. Related/concurrent works like [48] improve zero-shot reasoning by iteratively asking and answering questions with 3 VLM/LLM, and [30] uses scene-graphs to enhance compositional reasoning. [28, 51] devised sophisticated LLM-augmented tools for task subdivision and external tool selection. In contrast, this paper challenges the reasoning ability of VLMs on complex persuasive images through novel atypicality tasks and action-reason retrieval. It improves performance with a more lightweight atypicality-aware verbalization, and no external tools are needed.
Figure 2: Atypicality Understanding and Action-Reason Retrieval Tasks. We introduce three tasks: Multi-label Atypicality Classification , Atypicality Statement Retrieval, and Atypical Object Retrieval. Incorrect/correct phrases/statements are in red/green.
VLM evaluation. Various vision-language benchmarks have been proposed, including tasks like recognition [16], captioning [10], commonsense reasoning [40, 6], VCR [50], and compositional reasoning [37]. However, these benchmarks feature straightforward scenes that do not challenge the model’s reasoning abilities beyond literal imagery. Recently, ROME [53] generated counter-intuitive images focusing on 5 primitive common sense types (e.g., color and size) to challenge models’ ability in object, attribute, and spatial relation recognition. Similarly, WHOOPS! [7] addressed this by creating 500 synthetic scenes by placing ‘normal’ objects in unusual contexts (e.g., a snowplow in a desert and Einstein holding a smartphone). Atypical advertising images differ from those in WHOOPS! and ROME in two important ways: First, they are real ads created by human designers to intentionally convey a particular message (e.g., cigar replacing a bullet, to highlight the dangers of smoking), rather than simply to be unusual. This requires models to detect atypicality and atypical objects and reason about their impact on the ad’s meaning, making ad image understanding a more realistic and challenging benchmark for evaluating VLMs’ reasoning abilities. Second, atypical ads feature more categories of atypicality than placing an object out of context or altering primitive attributes, as WHOOPS! and ROME do, respectively.
3 Methodology
This paper evaluates VLM/LLM understanding of atypical advertisements. We address two key questions: (1) Are current VLMs capable of reasoning about atypicality and understanding advertisements? (2) What is the impact of atypicality on understanding ad images?
Unlike prior works [17] that only classify images as typical or atypical, we propose three new tasks: Multi-label Atypicality Classification (MAC), Atypicality Statement Retrieval (ASR), and Atypical Object Recognition (AOR). MAC predicts multiple categories of atypicality in the image. ASR uses additional annotations to identify objects involved in the atypical portrayal (e.g., "The surface of the bottle mimics the texture of feather"). AOR evaluates VLMs’ visual reasoning by identifying primary and secondary objects in atypical relation.
Our analysis shows while VLMs initially struggle with MAC and ASR tasks, they can extract valuable insights about atypical aspects of images. Leveraging these insights, we develop an atypicality-aware image verbalization. To detect atypicality, we use the prompt UH: What is unusual about this image?. Atypicality adds depth to the content of the advertisement, complementing surface-level content like objects and scene descriptions. Thus, we combine $UH$ and surface-level content to construct the final verbalization. It is then passed to an LLM for the action-reason inference task. We elaborate on the tasks and pipeline below.
Figure 3: Atypicality definitions and atypicality relation statements.
3.1 Proposed Atypicality Understanding Tasks
Ye et al. [46] devised a taxonomy of atypicality based on object transformations. In this work, we focus on the subset of atypicality categories that entail two objects (examples in Fig 1): (1) Texture Replacement 1 (TR1): Objects’ texture borrowed from another object, (2) Texture Replacement 2 (TR2): Texture created by combining several small objects, Object Inside Object (OIO): An object is completely or partially inside of another object, and (4) Object Replacement (OR): The whole object appearing in the context normally associated with another. We define the following new atypicality understanding tasks, shown in Fig. 2.
Multi-label Atypicality Classification (MAC). Unlike prior works [17] that only detect the presence of atypicality, we formulate atypicality detection as a multi-label classification task. The PittAds dataset provides three annotations of atypicality per image from different annotators, which may vary by type. For example, Fig.1(c) can be classified as ‘Object Inside Object’ (Earth inside a cup sleeve) and ‘Object Replacement’ (Earth replaces coffee cup). Some annotators may even label it as ‘Not Atypical’ (NA), creating five possible labels. MAC challenges VLMs to predict the relevant atypicality categories for an image based on atypicality definitions denoted as $\mathcal{D}_{\mathcal{A}}$ (prompts in supp). Due to the complexity of differentiating between these categories, we extend the definitions provided by [46] as shown in Fig. 3: they not only distinguish different atypicality categories but also hint at how atypicality impacts the image’s interpretation (e.g., Fig. 1 is TR1, where the beer’s texture is replaced by a feather to advertise its lightness).
(a) Image Verbalization
(b) Atypicality Statement Detection
(c) Action-Reason Retrieval
Figure 4: Our approach consists of three steps: (a) Image verbalization: We first prompt LLaVA to obtain top-5 objects ($V$), scene-text ($T$), scene description $IN$, and unusualness $UH$. Then we combine all the information to obtain a uniform description $\mathcal{T}_{\mathcal{V}}$. (b) Atypicality Statement Detection: We utilize $V$ and atypicality statement templates $\mathcal{S}_{\mathcal{A}}$ to generate the options which are then used along with $IN$ to retrieve the atypicality statement $\hat{s}$. (c) Action-Reason Retrieval: We input $\hat{s}$ along with verbalization $\mathcal{T}_{\mathcal{V}}$ to retrieve action-reason.
Atypicality Statement Retrieval (ASR). ASR formulates atypicality inference as retrieving a statement describing relations between two objects. We represent atypicality $\mathcal{A}$ with templates $\mathcal{S_{\mathcal{A}}}$, defined in Fig. 3. A statement $s=(a,o^{p},o^{s})$ includes an atypicality type $a$, primary object $o^{p}$, and secondary object $o^{s}$ as described in Sec. 3.6 of [46]. In TR1 and TR2, $o^{p}$ is the object with a new texture, and $o^{s}$ is source of that texture. In OIO, $o^{p}$ is the object inside, and $o^{s}$ is the object outside. In OR, $o^{p}$ is the object placed in another’s context, and $o^{s}$ is the expected object.
Given an ad, ASR distinguishes the correct atypicality statement ${s}^{+}=(a^{+},o^{+p},o^{+s})$, from a set of negative statements $\{\bar{s_{i}}\}\in{S}^{-}$. We generate negatives as follows: (1) Wrong object: replacing $o^{+p}$ and $o^{+s}$ with objects from $K$ random images, producing $2K$ negatives (e.g., $\bar{s}_{1}=(a^{+},o^{p}_{1},o^{s}_{1})$ and $\bar{s}_{2}=(a^{+},o^{s}_{1},o^{p}_{1})$ where $o^{p}_{1}/o^{s}_{1}$ are from a random image); (2) Wrong atypicality relation: altering the relation with one not in the ground-truths (objects remain the same) to create up to 3 negatives. (3) Swapped primary/secondary objects: we create $(a^{+},o^{+s},o^{+p})$. Thus, ASR tests the model’s understanding of objects, their atypicality relation, and primary/secondary object roles. It bridges the gap between MAC and action-reason retrieval by detecting atypical statements and enhancing action-reason retrieval. We use $K=2$.
Atypical Object Recognition (AOR). To assess the VLMs’ fine-grained visual understanding, we introduce the task of recognizing the primary and secondary objects in an atypical image. Given an ad and the true atypicality, the model must generate the correct primary/secondary objects to complete the atypicality statement. AOR can be viewed as a fill-in-the-blank task (i.e., generative) based on the statement templates.
3.2 Proposed Approach
To explore the impact of atypicality in action-reason retrieval and compare the reasoning ability of VLMs and LLMs, we propose a three-steps in-context learning method : (i) Atypicality-aware image verbalization: Using LLaVA [27] and LLM, we generate a coherent verbalization $\mathcal{T}_{\mathcal{V}}$ sensitive to atypicality; (ii) Atypicality statement detection: We detect the atypicality statement $\hat{s}=(\hat{a},\hat{o}^{p},\hat{o}^{s})$; (iii) Action-reason retrieval: We combine steps (i) and (ii) to retrieve the final action-reason statement. Fig. 4 illustrates the method.
(i) Verbalize image in atypicality-aware manner. Each ad image $I=\left(V,T\right)$ is composed of visual content ($V$) (objects) and textual content ($T$) (scene-text). We obtain $V$ and $T$ by querying LLaVA for up to 5 objects and a list of text-scenes visible in the image. However, this information is insufficient to fully comprehend the image. Hence, we additionally prompt LLaVA with (1) $ImageNarration$ ($IN$): LLaVA’s responses when prompted Describe the image in detail. (2) $UnusualHighlighter$ ($UH$): LLaVA’s reponse when asked What is unusual about the image?. MAC results (Table 1) show that $UH$ effectively captures image unusualness (closely related to atypicality), while $IN$ provides scene and object information useful for retrieving atypicality statement $\hat{s}$. Thus, we construct the final verbalization $\mathcal{T}_{\mathcal{V}}$ by combining $UH$, $V$, $T$, and $IN$ using an LLM.
(ii) Detect atypicality statement. We construct all possible statements to predict the atypicality statement $\hat{s}$. Option Generator (teal module in Fig 4(b)) combines $V$ and $\mathcal{S}_{\mathcal{A}}$ to generate all possible statements ${S}_{I}$. Specifically, each object pair $(o_{i},o_{j})\in V$ is combined with all atypicality statement template $s\in\mathcal{S}_{\mathcal{A}}$ to create atypicality statements $(a_{k},o_{i},o_{j})$ and $(a_{k},o_{j},o_{i})$ for all $a_{k}\in\mathcal{A}$. We then pass these atypical statements ${S}_{I}$ along with verbalization $\hat{S}_{IN}$into the classifier to predict $\hat{s}$ (no ground-truth is used).
(iii) Retrieve action-reason statements. [18] provided three action-reason statements for each image, each offering different plausible reasons for the same action. Given these three plausible (i.e., positive) and many implausible (i.e., negative) action-reason statements, the task is to detect all plausible action-reason statements $\{\hat{ar}_{I}\}_{i=1}^{3}$. We proposed various verbalization strategies, including concatenation and LLM-based combinations (i.e., $\mathcal{T}_{\mathcal{V}}$) of $V,T,IN,UH$, as well as concatenation of $\hat{s}$ with $\mathcal{T}_{\mathcal{V}}$, to be utilized by an LLM for retrieving the final action-reason statement.
4 Experimental Setup
Datasets. PittAds [18] includes 64,832 ad images, with 3,928 annotated for atypicality, action-reason, and primary/secondary objects. For MAC, we use atypicality categories, while for AOR and ASR, we use primary/secondary objects along with our atypicality statement templates $\mathcal{S}_{\mathcal{A}}$ (Fig. 3) to generate ground truth for evaluation. We utilized the train set with 1,168 samples for the main results, as it includes at least one annotation of the atypicality categories studied and is larger than the test set. No training was performed. Ablation studies are reported on a smaller subset of 250 images due to the high cost.
Baselines. We use LLaVA (‘vicuna-13b-v1.5’) [27], InstructBLIP (‘vicuna-13b-v0’) [13], MiniGPT4 [56] (‘vicuna-13b-v0’), and CLIP [33] (‘ViT-L/14@336px’ following [19]), and InternVL-Chat-V1-1 [11] as VLM baselines (and InternVL2-8B, LLaVA 1.6 in supp). LLaVA is our multimodal component due to its GPT-4-informed instruction tuning, state-of-the-art reasoning performance, and promptability [27]. We evaluate GPT-4V on a limited 250 examples, constrained by cost. We report BLIP-2 [24] (‘blip2-flan-t5-xl’) only for AOR as it failed to produce meaningful output when asked to return multiple options, i.e., MAC and multi-ARR (detail in supp).
Our analysis spans recent public LLMs, such as ‘vicuna-13b-v1.5’ (Vicuna) [38], ‘InternLM2-5-7b-chat’ (InternLM; see supp), and leading commercial models like GPT-3.5/4. We chose Vicuna as it is used in all VLMs (LLaVA, MiniGPT4, InstructBLIP) and InternLM as InternVL2-8B’s LM. We also compare GPT-4 and GPT-4V. We introduce CLIP ($I+T$) as a zero-shot baseline aligned with KAFA [19] but avoid direct comparison with KAFA as it is not publicly available. To assess our atypicality method, we compare against $V+T$ (verb. baseline), which includes basic image information (up to 5 objects, scene text).
Metrics. We use Precision, Recall, and F1-score to evaluate MAC (additional metrics in supp). For AOR, we assess sentence similarity between $s^{+}$ and $\hat{s}$ using ‘all-mpnet-base-v2’ [36], a state-of-the-art sentence embedding method. Common text-matching and accuracy metrics aren’t suitable for AOR since it is a generative task, and annotations can vary widely (e.g., ‘beer,’ ‘glass of beer,’ ‘beer glass’) due to human inconsistencies and typos. We report accuracy (denoted ‘Acc’) for single statement retrieval tasks, where the model returns only one statement per query (i.e., ASR and Single ARR). Top-k Acc and unranked Precision@k are the metrics for the multi-option ARR, with $Precision@k=\frac{min(k,\textit{\# of relevant statements in top k predictions})}{k}$. Note top-3 acc and prec@1 are the same.
Hard Negative (HN) Generation. To measure ARR performance, [45] mined hard negatives from images within the same topic, while [19] expanded the negative options. These negatives often include irrelevant objects, allowing VLMs to easily disregard them by comparing objects. This hinders accurate measurement of models’ reasoning ability. In contrast, a concurrent work [4] used annotators to write implausible statements based on visible objects/texts. Our approach differs in three key ways: (1) it is LLM-based, image-agnostic, and scalable; (2) it generates a wider variety of negatives while focusing on semantics (e.g., altering actions, reasons, adjectives, or swapping objects not visible such as ‘lipstick’ instead of ‘lip balm’), and (3) they evaluate contrastive-based VLMs, whereas we focus on generative VLM/LLMs with stronger reasoning ability.
Specifically, for each ground-truth action-reason statement, we ask GPT-4 to generate a negative action-reason statement by (1) Action alter: changing the action to an unrelated or opposite action while preserving the reason; (2) Reason alter: changing the reason to an unrelated or opposite reason; (3) Adjective alter: negating or modifying adjectives to make the statement incorrect; (4) Object swap: substituting at least one object in the statement with an unrelated object; (5) Statement alter: generating a completely unrelated action-reason statement.
We validated our hard negatives by sampling 100 images and asking annotators to select correct statements (in supp).
Implementation. For GPT-4, GPT-3.5, and Vicuna temperature is set to 0. For LLaVA [26], BLIP-2 [24], InternVL models [11] and InternLM [8] we used the original setting. We applied 8-bit quantization for LLaVA, MiniGPT-4, Vicuna, InternVL and InternLM. All experiments were zero-shot and conducted on an NVIDIA 368 Quadro RTX A5000 GPU with 24 GB memory. Example prompts are in supp.
5 Results
The key goal is to benchmark VLMs and evaluate their understanding of persuasive ads. This section first presents results on VLMs/LLMs’ atypicality understanding tasks, forming the foundation for our atypicality-aware verbalization. Then, we explore whether atypicality can help ad understanding on Action-Reason Retrieval (ARR).
5.1 Atypicality Understanding Results
We assess opone-sour VLMs’ (LLaVA [26], InstructBLIP [13]) understanding of the atypicality and GPT-4V. Additionally, we evaluate the effectiveness of two prompting strategies for capturing atypicality: (1) $IN$ (Describe the image in detail.) and (2) $UH$ (What is unusual about the image?). These strategies are compared to the VLMs and the $V+T$ baseline across three LLMs (GPT-4, GPT-3.5, Vicuna). Table 1 summarizes the results on MAC and ASR, while AOR results are in Table 2.
MAC ASR
Classifier Method Precision Recall F1-score Acc
$\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
LLaVA [27] I 27.75 27.75 42.38 52.71 21.24 26.03 18.83
IN 25.12 31.40 42.44 53.04 25.06 31.32 20.90
UH 44.35 30.44 42.04 52.44 24.16 29.98 17.90
InstructBLIP [24] I 34.81 27.60 41.43 50.73 17.72 20.18 19.76
Vicuna [12] $V+T$ 36.70 30.64 41.73 45.78 32.52 31.66 14.30
$IN$ 37.71 32.04 43.70 45.91 34.51 32.09 23.29
$UH$ 39.41 33.33 36.05 42.88 27.35 30.36 14.74
GPT 3.5 $V+T$ 41.46 35.36 23.21 21.54 28.18 24.95 50.00
$IN$ 46.28 42.50 25.13 14.75 28.49 19.64 50.55
$UH$ 49.10 43.34 27.38 30.92 27.06 28.24 50.05
GPT 4 $V+T$ 40.38 35.95 22.56 6.69 22.66 10.99 52.44
$IN$ 54.78 53.40 27.19 13.64 30.58 20.91 57.70
$UH$ 53.49 51.01 29.15 28.89 34.62 33.05 56.89
Table 1: Atypicality Understanding Tasks (MAC & ASR) on Full-set. $UH$ is very effective on MAC. $IN$ slightly outperforms $UH$ on ASR; ASR also requires to identify the objects which may not be well-represented in $UH$. $\checkmark$/$\times$ denotes performance with/without No Atypicality (NA) class. Best result per LLM bolded.
VLMs struggle with direct atypicality inference. Table 1 shows VLMs consistently underperform verbalization approaches ($UH$ and $IN$) in both MAC (F1-score) and ASR. For instance, LLaVA shows high recall (52.71) but low precision (27.75) in the MAC task, indicating it over-predicts atypicalities. This trend is particularly evident in the OIO and TR2 categories, where LLaVA achieves perfect and near-perfect recall (1.0 and 0.79) but low precision (0.18 and 0.23). Conversely, recall for other categories doesn’t exceed 0.24, suggesting a bias towards OIO and TR2 predictions (category-wise metrics not shown in table). InstructBLIP exhibits similar issues. This suggests VLMs lack the reasoning ability to accurately infer atypicality.
V+T lacks sufficient context to fully understand the image. $V+T$ provides inadequate context for extracting atypicality, as evidenced by GPT-4’s low recall/F1 (precision/recall scores of 58.11/86.04 for NA) and low recall for atypical categories (6.00, 0.79, 6.10, and 13.86 for TR1, TR2, OIO, and OR; not shown in table). Smaller LLMs (Vicuna and GPT-3.5) perform better than GPT-4 with V+T, but their improvement is likely due to hallucination rather than extracting useful information. In contrast, $UH$ and $IN$ provide richer descriptions that better capture atypicality.
VLMs are effective for verbalization. In MAC, $UH$ emerges as the top-performing strategy, effectively extracting unusualness and atypicality from images. It significantly improves $V+T$ F1 score without NA (details in section 3.1) by 3.29 and 22.06 for GPT-3.5 and GPT-4, respectively, and is only slightly lower than Vicuna’s $V+T$. $UH$ also outperforms $IN$ overall. In ASR, $UH$ is more effective than $V+T$ and outperforms VLMs by a significant margin when used in GPT models (e.g., by 37.13 with GPT-4 compared to the strongest VLM). However, $IN$ slightly outperforms $UH$ on GPT-4/GPT-3.5 and significantly on Vicuna. These results indicate that identifying atypicality requires both image understanding and reasoning capabilities. ASR necessitates the model to identify the correct atypicality statement, which includes both objects and their relation (e.g., search bar completely replaces mouth in its usual context, assuming its function or position in Fig. 1 (a)). Thus, $IN$ may be a better verbalization as it provides more detailed information about the image, objects, and their relations, whereas $UH$ offers implicit information about the image’s unusualness. Therefore, we use $UH$ (best on MAC) and $IN$ (best on ASR) along with $V+T$ to create atypicality-aware verbalization. Inspired by ASR results we also adopt $IN$ for detecting atypicality statement $\hat{s}_{IN}$ which is combined with our verbalization for ARR.
Model Avg. similarity ($\hat{s}$ to $s^{+}$) % of scores
score   $>$ 0.7    $>$ 0.6    $>$ 0.5
BLIP2 [24]  0.45   8.77    19.78    35.43
InstructBLIP [13]  0.46   9.54    21.24    40.76
MiniGPT4 [56]  0.51   15.24    32.28    51.71
LLaVA [27] 0.59   31.41    51.35    65.16
GPT-4V $\dagger$ [1]  0.67   46.94    61.63    77.14
Table 2: AOR on Full-set. $\dagger$ on Small-set.
VLMs’ limitations in Atypical Object Recognition. Table 2 explores VLMs performance on the AOR task, which involves generating primary and secondary objects given only GT atypicality relation to complete its atypicality statement from Fig. 3. Scores above 0.7 indicate strong semantic overlap, while scores between 0.5 and 0.7 indicate moderate overlap. The results underscore current VLMs’ difficulty in reasoning and recognizing atypical objects. InstructBLIP and MiniGPT4 perform poorly, with most predictions scoring below 0.5, highlighting their struggle to recognize primary and secondary objects in atypical contexts. GPT-4V emerges as the most proficient model, yet only about half of its predictions surpass the 0.7 mark. These results highlight the need to improve VLMs’ reasoning in complex visual scenes, such as atypical objects.
5.2 Action-Reason Retrieval Results
We evaluate our proposed method using four LLMs (GPT-4, GPT-3.5, Vicuna, InternLM in supp) and compare them against VLMs (CLIP, LLaVA, InstructBLIP, InternVL-Chat-V1-1, GPT-4V, LLaVA1.6, and InternVL2-8B (in supp)). Table 3 presents the multi-option action-reason retrieval results (ARR) and their comparison with VLMs on the full-set. Table 4 exhibits ARR results on small-set when GPT-4V is the VLM in our pipeline. We adopted Vicuna as the public LLM because it’s used in LLaVA, InstructBLIP, and MiniGPT4. We use GPT-4 and GPT-4V as powerful LLM/VLM pairs.
Classifier Verb. Precision@k Top-k Acc Avg
k=1 k=2 k=3 k=1 k=2
CLIP[33] $I$ 61.04 33.86 22.66 23.72 44.61 37.18
$I+T$ 46.15 24.36 16.24 15.15 31.25 26.63
LLaVA [27] $I$ 59.67 38.27 26.06 32.92 48.14 41.01
$I+\mathcal{T}_{\mathcal{V}}$ 59.45 37.37 25.14 27.49 47.07 39.30
InstructBLIP [13] $I$ 15.05 10.03 7.80 13.04 13.04 11.79
InternVL-Chat-V1-1 [11] $I$ 52.22 32.79 22.17 22.51 40.66 30.07
Vicuna [12] $V+T$ 64.13 40.71 27.57 21.49 43.41 39.46
$\mathcal{T}_{\mathcal{V}}$ (Ours) 67.38 44.01 29.94 23.20 41.95 41.30
$\mathcal{T}_{\mathcal{V}}+\hat{s}_{IN}$ (Ours) 68.32 44.52 30.25 22.95 43.24 41.86
$\mathcal{T}_{\mathcal{V}}$ (GPT-4 Verb.) (Ours) 68.49 44.52 30.37 24.06 43.24 42.14
GPT-4 $V+T$ 93.73 84.42 70.50 71.50 89.87 82.00
$\mathcal{T}_{\mathcal{V}}$ (Ours) 93.99 86.35 72.96 74.94 91.16 83.88
$\mathcal{T}_{\mathcal{V}}+\hat{s}_{IN}$ (Ours) 95.54 87.55 74.62 88.42 93.40 87.91
Table 3: ARR on Full-set . Predicted atypicality statement $\hat{s}$ uses the respective prompting (IN) and LLM (Vicuna, GPT-4), with all LLMs using LLaVA verbalization. $\mathcal{T}_{\mathcal{V}}$ combines $V+T+IN+UH$ using Vicuna or GPT-4. Best number per block and column is bolded. Precision@k is unranked. Task is Multi-ARR.
LLMs are more powerful than VLMs. In Table 3 and 4, LLMs with atypicality-aware verbalization ($\mathcal{T}_{\mathcal{V}}$) consistently outperform VLMs. For example, GPT-4 surpasses LLaVA by 42.87 points. GPT-4 also outperforms a strong VLM, GPT-4V, in all metrics (Table 4). Results in supp confirm this trend across InternVL2-8B and LLaVA 1.6. This highlights the superior reasoning ability of LLMs in understanding atypicality and action-reason statements.
Interestingly, when LLaVA is provided with both image and $\mathcal{T}_{\mathcal{V}}$ (LLaVA($I+\mathcal{T}_{\mathcal{V}}$)), it underperforms Vicuna by 7.93, 6.64, and 4.8 points on prec@1, prec@2, and prec@3, respectively. Also, its performance drops by 1.71 points compared to LLaVA($I$). This reveals that despite using Vicuna as its LLM, LLaVA’s reasoning ability is limited, hindering it effective use of $\mathcal{T}_{\mathcal{V}}$ for action-reason retrieval.
Figure 5: ARR error analysis on Full-set.
Classifier precision@k Top-k acc
k=1 k=2 k=3 k=1 k=2
LLaVA 59.67 38.27 26.06 32.92 48.14
GPT-4V 97.17 89.91 74.86 77.01 90.32
GPT-4 ($\mathcal{T}_{\mathcal{V}}$) 97.58 90.72 76.61 81.04 92.74
Table 4: GPT-4V verb. in Multi-ARR on Small-set.
Atypicality boosts persuasive visual media understanding. Table 3 demonstrates the effectiveness of our atypicality-aware verbalization method compared to VLMs and the $V+T$ verbalization baseline. While Vicuna($V+T$) lags behind LLaVA by 1.55 points on avg, indicating insufficient context from basic verbalization, our $\mathcal{T}_{\mathcal{V}}$ consistently improves $V+T$ performance. We observe gains of 3.8/2.8 and 1.93/2.46 on prec@2/prec@3 for Vicuna and GPT-4, respectively. Combining $\mathcal{T}_{\mathcal{V}}$ with $\hat{s}_{IN}$ yields the best results across all LLMs (GPT3.5 in supp), confirming the benefit of incorporating atypicality-aware verbalization and atypicality statement for understanding persuasive ads. Interestingly, Vicuna shows minimal gains when using GPT-4’s verbalization, suggesting that our strategy does not depend heavily on extensive LLMs like GPT-4.
Table 5 investigates how the addition of atypicality statements impacts VLM performance on ARR using three methods: (1) $I+s^{+}$, with true atypicality statements $s^{+}=(a^{+},o^{+p},o^{+s})$; (2) $I+\hat{s}$, with predicted atypicality statements $\hat{s}=(\hat{a},\hat{o}^{p},\hat{o}^{s})$ by GPT-4 and $\mathcal{T}_{\mathcal{V}}$; and (3) $I+\bar{s}$, with incorrect atypicality statements using correct objects but incorrect relations, $\bar{s}=(\bar{a},o^{+p},o^{+s})$. The results reveal that VLMs benefit from atypicality statements. However, for GPT-4 (an LLM), adding the atypicality to $\mathcal{T}_{\mathcal{V}}$ is not as useful, since $\mathcal{T}_{\mathcal{V}}$ already contains correct atypicality information, and incorrect statement reduce the performance. These findings highlight the importance of incorporating atypicality to improve VLM performance on ARR tasks.
Generalization to typical images. We compared our pipeline with Vicuna ($\mathcal{T}_{\mathcal{V}}$) on typical images (i.e., no atypicality) against LLaVA on small-set ARR (details in supp). Vicuna ($\mathcal{T}_{\mathcal{V}}$) achieved 71.2%/48.6%/33.2% vs. 66.4%/42.2%/28.3% for LLaVA on prec@1/2/3, demonstrating our approach’s effectiveness beyond atypical ads.
Figure 6: ASR error analysis on Full-set
Model $I$ $I+s^{+}$ $I+\hat{s}$ $I+\bar{s}$
LLaVA 26.00 35.18 54.28 28.16
InstructBLIP 20.44 23.25 23.40 19.69
GPT-4V 86.87 89.35 87.24 86.96
GPT-4 ($\mathcal{T}_{\mathcal{V}}$) 96.77 91.42 96.76 90.20
Table 5: Atypicality ablation on ARR Small-set. $I$: image, $s^{+}$: GT atyp., $\hat{s}$: predicted atyp. based on $\mathcal{T}_{\mathcal{V}}$, $\bar{s}$: incorrect atyp. Acc on Single ARR. Best result per row in bold.
5.3 Further Analysis & Ablation
Neg. Strategy Model Multi Single
Precision@k
k=1 k=2 k=3 Acc
12 Neg. [18, 19] CLIP $(I)$ 98.79 97.58 92.20 96.77
CLIP $(I+T)$ 97.58 97.58 87.10 90.32
LLaVA $(I)$ 93.47 74.08 56.33 94.31
GPT4 $(\mathcal{T}_{\mathcal{V}})$ 99.60 96.98 91.13 93.52
18 Hard Neg. CLIP $(I)$ 64.52 34.48 22.98 20.97
CLIP (I+T) 47.18 25.40 16.94 15.73
LLaVA $(I)$ 59.67 38.27 26.06 26.80
GPT4 $(\mathcal{T}_{\mathcal{V}})$ 96.77 87.30 74.60 96.77
Table 6: Impact of semantically hard negatives on Small-set.
Hard Negatives ablation. Table 6 compares our semantically hard negatives against those used in prior work. VLM performance drops substantially when faced with our hard negatives, evidenced by 69.22/30.27 drop in CLIP(I)/LLaVA(I) on prec@3. Conversely, our method exhibits a decrease of no more than 17 across all metrics, demonstrating robustness in reasoning compared to VLMs.
Error analysis on ARR. In Fig. 5, VLMs (i.e., LLaVA and CLIP) do not make considerably more errors than GPT-4 ($\mathcal{T}_{\mathcal{V}}$) on ‘object swap’ since these negative options contain different objects from the ground-truth, making them easier for VLMs to identify. However, VLMs make substantially more errors than GPT-4($\mathcal{T}_{\mathcal{V}}$) on semantically incorrect negatives (i.e., ‘action alter,’ ‘reason alter,’ ‘statement alter,’ and ‘adjective alter’). This confirms VLMs mainly rely on visual elements (e.g., objects) rather than deeper reasoning (examples in supp).
Error analysis on ASR. In Fig. 6, LLaVA makes notably more errors, particularly on ‘Wrong Relation’ options, which demand deeper reasoning than other negative types.
Effectiveness of each component in atypicality-aware verbalization. We ablate the effectiveness of each step in atypicality-aware verbalization on the ARR small-set (details in supp). The performance of $\mathcal{T}_{\mathcal{V}}$ shows the advantage of atypicality-aware verbalization over basic concatenation ($V+T+IN+UH$). Specifically, $\mathcal{T}_{\mathcal{V}}$ improves top-1 acc by 14.76 on GPT-3.5 and 12.78 on GPT-4. This is because LLaVA-generated descriptions are inherently noisy, and their naive concatenation can mislead the models. A further issue is increased prompt length. Additionally, $UH$ alone is less effective than $IN$, but $\mathcal{T}_{\mathcal{V}}$ outperforms $\mathcal{T}_{\mathcal{V}}\setminus UH$, showing that atypicality is important yet complementary to basic image information captured in $V$, $T$, and $IN$.
Generalization beyond ads. We test our atypicality-aware verbalization pipeline on WHOOPS! [7] in supp.
6 Conclusion
This work challenged VLMs on complex rhetorical visual media, focusing on atypicality in advertisements. We introduced three novel atypicality tasks and benchmarked VLMs on and ARR, revealing their limitations in advanced reasoning. Despite these limitations, VLMs showed potential in extracting relevant information for understanding the atypical images. Our atypicality-aware verbalization strategy significantly enhances LLM performance on ARR tasks. Extensive experiments demonstrate that our approach outperforms VLM baselines, proving the effectiveness of incorporating atypicality inference for understanding ad images. These findings highlight the importance of atypicality in interpreting complex visual media and the superior reasoning abilities of LLMs over VLMs.
Limitations. The PittAd dataset [18] is widely used for understanding visual media (e.g., [45, 2, 19, 17, 31]), but it contains many annotations reflecting human biases and some images with sensitive content. Exploring these biases is beyond the scope of this paper.